This is a homework problem due tomorrow. 4 people have done it and we have 4 seperate answers. Help!
Solve the equation.
t^2 + 1 + (13/6)t

Answer by checkley71(8403)

T~2+13/6T+1=0 OR (x+2/3)(x+1.5)=0 or x=-2/3 & x=-1.5
I FORGOT TO ADD THE FOLLOWING STEPS IN THE SOLUTION X+2/3=0 OR X=-2/3 &
X+1.5=0 OR X=-1.5
SORRY!!!!